Consider a sample with data values of 27, 25, 20, 15, 30, 34, 28, and 25. Compute the 22nd, 27th, 64th, and 74th percentiles. If needed, round your answers to two decimal digits.
Answer #1

Xth Percentile =(x*n)/100 value

27, 25, 20, 15, 30, 34, 28, and 25.

We will arrange above data in ascending order

15,20,25,25,27,28,30,34

n=8

22nd percentile=(22*8)/100 =1.76 i.e. second value which is 20

27th percentile=(27*8)/100 =2.16 i.e. Third value which is 25

64th percentile=(64*8)/100 =5.12 i.e. sixth value which is 28

74th percentile=(74*8)/100 =5.92 i.e. sixth value which is 28